OpenStudy (anonymous):

Solve and graph the absolute value inequality: |2x + 1| ≤ 5.

OpenStudy (texaschic101):

| 2x + 1 | <= 5 -5 <= 2x + 1 <= 5 --- subtract 1 -5 - 1 <= 2x <= 5 - 1 -6 <= 2x <= 4 --- divide by 2 -3 <= x <= 2 so we have x being greater or equal to -3 and less then or equal to 2 closed circle on -3, closed circle on 2, shading in between. Try to learn this : if there is an equal sign, there is a closed circle no equal sign means an open circle < less then means shading to the left > greater then means shading to the right or just shade the way the inequality sign is pointing.

OpenStudy (texaschic101):

you can read up on this here... http://www.purplemath.com/modules/absineq.htm

OpenStudy (anonymous):

thankyou i have a question can i message you ? @texaschic101

OpenStudy (anonymous):

i choose A

OpenStudy (texaschic101):

look my answer again and find the graph that shows x is greater or equal to -3 and less then or equal to 2. Now what do you think the answer is now...it is not A

OpenStudy (texaschic101):

ok....I am in a good mood...answer is B

OpenStudy (anonymous):

hmmm ok thanks I'm going to try too see where i went wrong I'm a little confused @texaschic101

OpenStudy (anonymous):

i understand now the link really worked thank you